# Online Difficulty Filtering for Reasoning Oriented Reinforcement Learning

### v1

@@ Appendix B Learnability in Soft Prompts
$\mathbb{D}_{\mathrm{KL}}\left(\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})\|\pi^{*}(\mathbf{y}|\mathbf{x})\right)\geq\frac{p(\mathbf{x})(1-p(\mathbf{x}))}{2\beta^{2}},$ (34)
explicitly establishing the Bernoulli variance scaled by $\frac{1}{2\beta^{2}}$ as a lower bound of the KL divergence between the initial policy and the optimal policy.
Appendix C Training Configurations
All experiments are built on the Qwen2.5-3B base model (Yang et al., 2024a). We integrate DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) optimization in our training pipeline to handle memory and computation efficiently. Both the SFT and RORL stages are conducted on a distributed setup of 8$\times$×NVIDIA A100 (80GB) GPUs.
@@ Appendix C Training Configurations
We used a learning rate of $5\times 10^{-6}$ and fine-tuned it for 5 epochs. The learning rate schedule was linear, with the first 25 steps used for warm-up. We used a batch size of 21.
@@
To guide the model toward producing responses aligned with the DeepSeek R1 format, we introduce a format reward based on five constraints: (1) the response must begin with a ‘$<$think$>$’ tag, (2) the ‘$<$think$>$’ section must be properly closed with a ‘$<$/think$>$’ tag, (3) the ‘$<$think$>$’ section must be non-empty, (4) the summary section following ‘$<$/think$>$’ must also be non-empty, and (5) the response must terminate with an eot token. Each constraint contributes 0.2 points, resulting in a maximum format reward of 1.0. In addition, we implement a language reward to reduce language mixing, especially given that all prompts during training and evaluation are in English. This reward was computed as the ratio of characters in the response that are alphabetic, symbolic (e.g., mathematical symbols), or whitespace, and ranged from 0 to 1. Lastly, we define an accuracy reward, assigning a score of 1.0 for correct answers and 0.0 for incorrect ones. The total reward is the sum of these three components—format, language, and accuracy—yielding a final reward score between 0 and 3.
Appendix D Evaluation Benchmarks
@@
  • OlympiadBench (He et al., 2024) includes 674 open-ended text-only competition problems from a broader set of 8,476 Olympiad and entrance exam questions, specifically using the OE_TO_maths_en_COMP subset.

### v2

@@ Appendix C Learnability in Soft Prompts
$\mathbb{E}\!\left[\log\frac{\pi^{*}(y\mid x)}{\pi_{\mathrm{init}}(y\mid x)}\right]\;\leq\;-\frac{p(x)\big(1-p(x)\big)}{2\beta^{2}}+\mathcal{O}\!\left(\frac{1}{\beta^{3}}\right).$ (33)
By the reverse KL identity in equation 9,
$D_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}(\cdot\mid x)\,\|\,\pi^{*}(\cdot\mid x)\big)\;=\;-\mathbb{E}\!\left[\log\frac{\pi^{*}(y\mid x)}{\pi_{\mathrm{init}}(y\mid x)}\right]\;\geq\;\frac{p(x)\big(1-p(x)\big)}{2\beta^{2}}+\mathcal{O}\!\left(\frac{1}{\beta^{3}}\right).$ (34)
Dropping higher-order terms establishes the stated bound:
$D_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}(\cdot\mid x)\,\|\,\pi^{*}(\cdot\mid x)\big)\;\geq\;\frac{p(x)\big(1-p(x)\big)}{2\beta^{2}}.$ (35)
Extremal cases.
If $p(x)\in\{0,1\}$, then $r_{\mathrm{acc}}(x,y)$ is almost surely constant under $\pi_{\mathrm{init}}$, so $V^{*}(x)=\mathbb{E}[r_{\mathrm{acc}}(x,y)]$ and equation 30 evaluates to 0. Hence
$D_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}(\cdot\mid x)\,\|\,\pi^{*}(\cdot\mid x)\big)=0,$ (36)
which matches Proposition 3.1 at the endpoints.
$\square$
Appendix D Training Configurations
[…]
Training Data Curation
For SFT, we sample problems from the NuminaMath dataset (Li et al., 2024) and generate solutions using DeepSeek-R1 (Guo et al., 2025). Only samples with verifiably correct solutions are retained, and we stop once approximately 1,000 such problem-solution pairs are collected. The final SFT dataset contains 1,107 filtered problems. For RLVR, we adopt a subset of the public dataset used in Cui et al. (2025) [^1] https://huggingface.co/datasets/PRIME-RL/Eurus-2-RL-Data. We specifically use only the math domain problems. This dataset provides a diverse pool of challenging prompts.
@@ Appendix D Training Configurations
We use a learning rate of $5\times 10^{-6}$ and fine-tune it for 5 epochs. The learning rate schedule is linear, with the first 25 steps used for warm-up. We use a batch size of 21.
@@
To guide the model toward producing responses aligned with the DeepSeek R1 format, we introduce a format reward based on five constraints: (1) the response must begin with a ‘<<think>>’ tag, (2) the ‘<<think>>’ section must be properly closed with a ‘<</think>>’ tag, (3) the ‘<<think>>’ section must be non-empty, (4) the summary section following ‘<</think>>’ must also be non-empty, and (5) the response must terminate with an eot token. Each constraint contributes 0.2 points, resulting in a maximum format reward of 1.0. In addition, we implement a language reward to reduce language mixing, especially given that all prompts during training and evaluation are in English. This reward was computed as the ratio of characters in the response that are alphabetic, symbolic (e.g., mathematical symbols), or whitespace, and ranged from 0 to 1. Lastly, we define an accuracy reward, assigning a score of 1.0 for correct answers and 0.0 for incorrect ones. The total reward is the sum of these three components—format, language, and accuracy—yielding a final reward score between 0 and 3.
Appendix E Evaluation Benchmarks
@@
  • OlympiadBench (He et al., 2024) includes 674 open-ended text-only competition problems from a broader set of 8,476 Olympiad and entrance exam questions, specifically using the “OE_TO_maths_en_COMP” subset.
@@
Appendix F Difficulty-Aware Performance Analysis
To further understand the effect of our method, we analyze performance variations based on difficulty levels.
Benchmark-Level Difficulty Spectrum
As discussed in § 4, our benchmark suite spans a wide difficulty range. This is reflected in the SFT checkpoint performance of Qwen2.5-3B, which ranges from 0.0% to 49.8% as shown in Table 1. We order the benchmarks in ascending difficulty according to SFT performance: AIME (0.0%), MinervaMath (13.2%), OlympiadBench (17.3%), AMC (20.5%), and MATH500 (49.8%). From this perspective, we observe two trends:
  • Narrowing the difficulty threshold (i.e., tighter filtering range) generally improves performance, especially on challenging tasks like MinervaMath and AIME.
  • Harder benchmarks benefit more from filtering. For instance, AIME shows more than a 300% relative improvement over SFT, and MinervaMath improves by 35%.
Difficulty-Level Breakdown within MATH500
We also analyze performance by difficulty levels in the MATH500 benchmark. Table 3 shows that balanced filtering GRPO outperforms plain GRPO across most difficulty levels, especially on harder ones (Level 3–5).
Difficulty Plain ($0\leq p(\mathbf{x})\leq 1$ ) w/ Online Filtering ($0<p(\mathbf{x})<1$) w/ Online Filtering ($0.3<p(\mathbf{x})<0.7$)
Level 1 88.37 88.37 83.72
Level 2 78.89 83.33 83.33
Level 3 71.43 70.48 79.05
Level 4 47.66 50.78 55.47
Level 5 30.60 32.84 32.09
Table 3: Accuracy (%) of GRPO-trained models on MATH500 by difficulty level. The highest score for each level is in bold.
Appendix G General rewards: CGF derivations and bounds
We provide full proofs for Proposition 6.1, Corollary 6.2, and Corollary 6.3.
G.1 Proof of Proposition 6.1.
Proof. Fix $x$ and write expectations over $y\sim\pi_{\mathrm{init}}(\cdot|x)$. By the definition of the soft value,
$V^{*}(x)\;=\;\beta\,\log\mathbb{E}\!\left[\exp\!\left(\tfrac{1}{\beta}\,r(x,y)\right)\right].$ (37)
From equation 9,
$\mathbb{E}\!\left[\log\frac{\pi^{*}(y|x)}{\pi_{\mathrm{init}}(y|x)}\right]\;=\;\frac{1}{\beta}\,\mathbb{E}[r(x,y)]\;-\;\frac{1}{\beta}\,V^{*}(x)\;=\;\frac{1}{\beta}\,\mu(x)\;-\;\log\mathbb{E}\!\left[e^{r/\beta}\right].$ (38)
Multiplying both sides by 1-1 gives
$D_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}\|\pi^{*}\big)\;=\;-\,\mathbb{E}\!\left[\log\frac{\pi^{*}}{\pi_{\mathrm{init}}}\right]\;=\;\log\mathbb{E}\!\left[e^{r/\beta}\right]\;-\;\frac{1}{\beta}\,\mu(x).$ (39)
Factor out the mean inside the exponential:
$\log\mathbb{E}\!\left[e^{r/\beta}\right]\;=\;\log\mathbb{E}\!\left[e^{(r-\mu)/\beta}\,e^{\mu/\beta}\right]\;=\;\frac{\mu}{\beta}\;+\;\log\mathbb{E}\!\left[e^{(r-\mu)/\beta}\right].$ (40)
Substitute into the previous line to obtain
$D_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}\|\pi^{*}\big)\;=\;\log\mathbb{E}\!\left[\exp\!\left(\tfrac{r-\mu}{\beta}\right)\right]\;=\;K_{\,r-\mu}\!\left(\tfrac{1}{\beta}\right),$ (41)
which is the desired identity. $\square$
G.2 Proof of Corollary 6.2.
Proof. Let $Z:=r_{G}-\mu_{G}$. Under $\pi_{\mathrm{init}}$, $Z\sim\mathcal{N}(0,\sigma_{G}^{2})$, whose moment generating function is $\mathbb{E}[e^{tZ}]=\exp(\tfrac{1}{2}\sigma_{G}^{2}t^{2})$. Therefore the CGF is $K_{Z}(t)=\tfrac{1}{2}\sigma_{G}^{2}t^{2}$. Applying Proposition 6.1 with $t=\tfrac{1}{\beta}$ yields
$D_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}\|\pi^{*}\big)\;=\;K_{Z}\!\left(\tfrac{1}{\beta}\right)\;=\;\frac{\sigma_{G}^{2}}{2\,\beta^{2}}.$ (42)
$\square$
G.3 Proof of Corollary 6.3.
Proof. Let $Z:=r_{M}-\mu$. Write the CGF Taylor series around $t=0$:
$K_{Z}(t)\;=\;\sum_{k=2}^{\infty}\frac{\kappa_{k}}{k!}\,t^{k}\;=\;\frac{\kappa_{2}}{2}\,t^{2}\;+\;\frac{\kappa_{3}}{6}\,t^{3}\;+\;\frac{\kappa_{4}}{24}\,t^{4}\;+\;\cdots,$ (43)
where $\kappa_{2}=\sigma^{2}=\mathrm{Var}(Z)$ and $\kappa_{k}$ are the cumulants of $Z$. Because $r_{M}$ is bounded (multinomial or $[a,b]$), all moments of $Z$ are finite and thus all $\kappa_{k}$ are finite. For $t=\tfrac{1}{\beta}$,
$K_{Z}\!\left(\tfrac{1}{\beta}\right)\;=\;\frac{\sigma^{2}}{2\,\beta^{2}}\;+\;\frac{\kappa_{3}}{6\,\beta^{3}}\;+\;\frac{\kappa_{4}}{24\,\beta^{4}}\;+\;\cdots.$ (44)
Hence,
$K_{Z}\!\left(\tfrac{1}{\beta}\right)\;\geq\;\frac{\sigma^{2}}{2\,\beta^{2}}\;-\;\left|\frac{\kappa_{3}}{6\,\beta^{3}}\right|\;-\;\left|\frac{\kappa_{4}}{24\,\beta^{4}}\right|\;-\;\cdots,$ (45)
which shows that the variance term dominates at second order and provides a lower control up to $\mathcal{O}(\beta^{-3})$. In particular, for sufficiently large $\beta$ (or small $t$), we obtain
$K_{Z}\!\left(\tfrac{1}{\beta}\right)\;\geq\;\frac{\sigma^{2}}{2\,\beta^{2}}\;-\;\mathcal{O}\!\left(\frac{1}{\beta^{3}}\right).$ (46)
Finally, apply Proposition 6.1 to translate this bound to the reverse KL, completing the proof. $\square$
Appendix H Discussion: Native Generalizability of Online Difficulty Filtering in RL for LLMs
The concept of online difficulty filtering in reinforcement learning for language models is inherently versatile across verifiable tasks and reward formulations, even beyond the empirical scope of this paper. Recent work in reinforcement learning with verifiable rewards (RLVR) demonstrates its applicability to diverse domains, including logic puzzles (Chen et al., 2025a), medical reasoning (Gunjal et al., 2025), complex instruction-following (Pyatkin et al., 2025), and coding (Chen et al., 2025b). Our theoretical framework builds on the standard KL-regularized or soft reinforcement learning formulation (Schulman et al., 2018; Richemond et al., 2024), which optimizes the objective
$\max_{\pi_{\theta}}\;\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}[\,r(x,y)\,]-\beta\,\mathrm{D}_{\mathrm{KL}}\!\big(\pi_{\theta}(\cdot|x)\,\|\,\pi_{\mathrm{init}}(\cdot|x)\big),$ (47)
yielding the Boltzmann-rational optimal policy relative to the reference $\pi_{\mathrm{init}}$. Modern post-training algorithms such as PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), RLOO, and REINFORCE++++ (Hu et al., 2025) can all be interpreted as optimizing surrogates of this KL-regularized objective, differing primarily in how the policy gradient is estimated or regularized. Hence, the theoretical mechanism behind online difficulty filtering, namely that the reverse KL divergence between $\pi_{\mathrm{init}}$ and $\pi^{*}$ is determined by the cumulant generating function of the centered reward and dominated by its variance, extends directly to these algorithms as long as a KL penalty to the reference is preserved.
The generalized analysis in §6.3 further shows that this relationship holds exactly for Gaussian rewards and forms a tight second-order lower bound for multinomial or bounded rewards. Since continuous neural reward models are approximately Gaussian and rubric-based verifiable rewards are discrete and bounded, these two cases jointly cover most practical reward systems in RLHF and RLVR. Accordingly, the same variance-guided filtering principle applies regardless of task or reward type: prompts with intermediate success probabilities maximize the effective learning signal, while trivially easy or hard ones contribute negligible gradients. This makes balanced online difficulty filtering a theoretically grounded and natively general component for PPO-, GRPO-, or RLOO-style alignment training across diverse verifiable tasks.